Circular No. 6678

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA SCORPII 1997 M. Cavagna, Sormano, Italy, reports the following precise position of

Liller’s possible nova (IAUC 6675), from a CCD image obtained on June 7.011 UT (0.5-m reflector): α = 17^{h}54^{m}11^{s}.22, δ = −30^{o}02′53′′.0 (equinox 2000.0; estimated mag 9). G. J. Garradd, Loomberah, N.S.W., reports the following position end figures, from a CCD image obtained on June 7.511 (0.25-m f/4.1 reflector): 11^{s}.19, 52′′.9 (mag 9.2). The Digital Sky Survey shows no definite object at this location, though a USNO A1.0 star of mag R = 18.7, B = 19.0 has position end figures 11^{s}.53, 54′′.3.

M. Fujii, Okayama; and H. Kawakita, Osaka, Japan, report that spectroscopic observations (resolution ∼ 1.2 nm; slit size 424′′ × 7′′) on June 7.62 UT show strong, broad Hα emission (FWHM ∼ 3200 km/s; 0.28-m reflector, instrumental resolution ∼ 550 km/s).

A. C. Gilmore, Mount John University Observatory, reports the following photometry of Liller’s variable star (which is the middle of three stars of similar brightness that are ∼ 1′ apart, aligned east-west), obtained with the 0.6-m f/16 reflector: June 7.507 UT, V = 10.48 ± 0.01, U − B = −0.17 ± 0.03, B − V = +0.89 ± 0.02, V − R = +1.31 ± 0.01, V − I = +1.95 ± 0.01 (comparison stars E790 and E782 from Cousins’ E region).

Visual magnitude estimates: June 7.15 UT, 10.1 (A. Pereira, Cabo da Roca, Portugal); 7.99, 10.4 (M. Plsek, Lelekovice, Czech Republic); 7.991, 9.8 (K. Hornoch, Lelekovice, Czech Republic); 8.12, 10.9 (Pereira).

AX J1820.5–1434 K. Torii, K. Kinugasa, and S. Kitamoto, Osaka University, report:

“A new x-ray pulsar, AX J1820.5–1434, has been discovered with ASCA during a galactic-plane survey project on Apr. 9 at α = 18^{h}20^{m}29^{s}.5, δ = −14^{o}34′24′′ (equinox 2000.0; preliminary error radius 0′.5). A coherent pulsation was detected from the source with an apparent barycentric pulse period of 152.23 ± 0.05 s. The mean flux, not corrected for interstellar absorption, was 2.4 × 10^{−11} erg cm^{−2} s^{−1} in the energy range 1–10 keV. The 1–10-keV pulse shape has one broad peak with a phase width of 0.3. At the peak, the pulsed flux is comparable with the flux of the constant component. The energy spectra obtained by both the GIS and the SIS were fitted by a power-law model (photon index of 1.0 ± 0.4) with a high column density of (1.0 ± 0.3) × 10^{23} H atoms cm^{−2}.”

1997 June 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green